# Euclid Quick Data Release (Q1)

### v1

@@ 6.3 LEMON applied to real Euclid Q1 lenses
We have finally applied LEMON to the sample of strong lens images found by the strong lensing discovery engine in the Euclid Quick Release 1 (Q1) data (Euclid Collaboration: Walmsley et al. 2025). Through an initial scan of the Q1 galaxy catalogues with machine-learning models, followed by citizen science inspection, expert vetting, and system-by-system modelling, Euclid Collaboration: Walmsley et al. (2025) has collected 500 strong lens candidates, divided in Grade A (246) and B (254) lens candidates after the expert vetting scoring. We complement these data with 38 Grade A and 40 Grade B candidates found after the expert visual classification performed on high-velocity dispersion galaxies in Euclid Collaboration: Rojas et al. (2025).
@@
The strong lenses in Euclid Collaboration: Walmsley et al. (2025); Euclid Collaboration: Rojas et al. (2025) have been modelled via PyAutoLens, which is the standard algorithm for lens modelling in Euclid  [^8] https://github.com/Jammy2211/PyAutoLens (Nightingale & Dye 2015; Nightingale et al. 2018, 2021b). PyAutoLens is an open-source Python package for strong gravitational lensing, which includes both fully automated strong lens modelling of galaxies and galaxy clusters, and tools for simulating samples of strong lenses. In this work, PyAutoLens employs an SIE mass model, consistent with the model used to train LEMON.
[…]
Figure 13: Same as Figs. 8 and 12, but for the Euclid lenses found in the Q1 fields.
Table 4: Same as Table 3, but for the Euclid lenses found in the Q1 fields.
Metrics $R_{\textrm{Ein}}/\textrm{arcsec}$
Bias ($\mu$) 0.05
Lower discr. scatter ($\sigma^{-}$) 0.09
Upper discr. scatter ($\sigma^{+}$) 0.22
RMSE 0.36
MAE 0.20
NMAD 0.14
Coefficient of determination ($R^{2}$) 0.49
7 Speed-up of traditional methods due to LEMON
@@
One possible use of LEMON in the context of strong lensing applications within Euclid, beside fast modelling of strong gravitational lenses, is to speed-up the modelling of strong gravitational lenses by giving traditional lens modelling methods initial starting points from which to start the search.
@@
To perform lens modelling, PyAutoLens adopts the Python library PyAutoFit [^9] https://github.com/rhayes777/PyAutoFit(Nightingale et al. 2021a). The standard modelling pipeline uses nested-sampling algorithms, such as nautilus (Lange 2023), for Bayesian inference of the lens model. To efficiently use the information from LEMON, PyAutoLens can use a gradient ascent optimiser, which fully exploits the information from the initial starting point (Amorisco et al. 2022). The lens light and source galaxy during lens modelling are modelled using a Multi-Gaussian expansion (MGE) technique (see He et al. 2024), independent of any information provided by LEMON, most notably meaning that, when using the gradient ascent optimiser, the centre of the source galaxy uses an initial starting point at (0,0)0″0″(\$\$,\$\$)( 0  ″ , 0  ″ ).
[…]
@@
Finally, to assess whether the optimiser approach needs LEMON in order to correctly model gravitational lenses, we repeated the test by using random starting points for the optimiser. Results show that, with random initialisations, PyAutoLens is unable to successfully recover a correct lens model fit for any of the lenses. Thus, the robust initial starting points provided by LEMON prove to be required in order for PyAutoLens to be sped up by using the gradient ascent optimiser instead of the nested sampling algorithm.
An important benefit of combining LEMON and PyAutoLens lies in their ability to complement each other’s limitations. For instance, the standard PyAutoLens modelling pipeline inferred an incorrect lens model for one system due to misinterpreting additional line-of-sight emission from an unrelated interloper galaxy as part of the lensed source. By using LEMON as a starting point, the initial solution was already aligned with the strong lens, allowing PyAutoLens to refine the model through gradient ascent without being misled by interloper emission. Running independent fits with both the standard PyAutoLens pipeline and the combined LEMON-PyAutoLens approach for the same lens, and selecting the best-fit model from the two therefore provides a viable strategy to minimise failure rates in lens modelling.
@@ 8 Conclusions
In this work, we expand the lens modelling machine-learning algorithm LEMON (Gentile et al. 2023) by predicting both the mass and the light of the foreground deflector of Euclid galaxy-galaxy gravitational lenses. The former is modelled as an SIE without external shear, while the latter is a single Sérsic profile. We used mock Euclid lenses with contaminants in the image for both training and testing of the architecture. The main results are as follows.
@@
  • All parameters of the test set are recovered correctly, with very low values of bias and scatter. The parameter that is recovered worst is the Sérsic index, which shows a large scatter and a deviation from the linear trend for high values of the Sérsic index. This suggests that LEMON is able to recover the mass and light parameters for the lens even in the presence of companions contaminating the image.
  • The epistemic uncertainties are systematically lower than the aleatoric uncertainties, which is consistent with a training set that correctly spanned the parameter space.
  • Applied to a selection of Euclidised HST images, LEMON manages to recover Einstein radius, axis ratio, position angle, and lens effective radius of the lens, while we cannot reach any conclusion on the Sérsic index due to low statistics. As expected, in general results are worse than those obtained from the test set, given that Euclidised lenses have a much worse S/N than the simulated lenses, and may not be similar to the simulations used to train the network, as shown in Fig. 10. It should also be noted that both SLACS and ACS lenses have been modelled by assuming a De Vaucouleurs light profile, while LEMON assumes a Sérsic profile with variable values for the Sérsic index, and as such the associated true values for the lens effective radius cannot be compared homogeneously with the LEMON predictions. Moreover, a number of lenses have effective radii bigger than $2\arcsec$, which is the upper limit of the respective training sample for LEMON, and as such their predictions are unreliable extrapolations, and have not been reported. In the future, we will expand the training sample to include lenses with effective radii bigger than $2\arcsec$.
@@
  • Applied to five real galaxy-galaxy Euclid gravitational lenses observed in the Perseus ERO field and the 578 new candidates in the Q1 data (Acevedo Barroso et al. 2024; Euclid Collaboration: Walmsley et al. 2025; Euclid Collaboration: Rojas et al. 2025), for which the value of the Einstein radius is available thanks to classical modelling, we obtain an accurate recovery of the parameter for all lenses.
@@
  • By using the predictions of LEMON as starting points for PyAutoLens, it is possible to speed up the algorithm by up to 26 times compared to the standard Euclid modelling pipeline. This would not be possible without the initial LEMON guesses, because the gradient-ascent optimiser with random initialisations does not produce valid model results for all lenses.
@@
In future works, we plan to further advance the LEMON architecture, by predicting the parameters of more general lens mass models, such as the power-law model. We also plan on enhancing the prediction of certain parameters by deblending the gravitational lens image in lens-only and source-only images (Zhong et al. 2024), and train LEMON separately on the two components. We will also test in more detail the recovery of the lens light parameters, predicting values of effective radius and Sérsic index of real Euclid data, when they will become available. Finally, we will try to upgrade LEMON to also predict the parameters of the source, especially its position, given its usefulness in enhancing the performance of the Euclid standard modelling pipeline.
@@
Q1 data are now public, with DR1 data around the corner. We are therefore preparing to process the enormous amount of upcoming data, and have started applying LEMON to newly discovered strong lenses from the Q1 data release. Results confirm that LEMON is a robust and fast lens modelling tool for Euclid. This will enable the generation of models for the approximately 100 000100000100\,000100 000 expected Euclid lenses in a remarkably short time.
@@
By applying LEMON and PyAutoLens to Euclid gravitational lenses, we will determine the total mass within the Einstein radius, or extrapolated to the effective radius, when lens and source redshifts are available. Combining this with stellar mass estimates from SED fitting (Euclid Collaboration: Enia et al. 2025), we will derive dark matter fractions. We will also provide constraints on the mass density slope and initial mass function. With such a vast statistical lens sample, we will be able to investigate the evolution of these properties in unprecedented detail, as well as the scaling relations which relate them as a function of redshift (Koopmans et al. 2006; Gavazzi et al. 2007; Bolton et al. 2008; Auger et al. 2010; Tortora et al. 2010, 2014, 2018, 2019; Sonnenfeld et al. 2013, 2015, 2017; O’Riordan et al. 2025). By comparing these scaling relations with predictions from cosmological simulations, we will be able to constrain cosmological parameters and physical processes in galaxies (Mukherjee et al. 2018, 2021, 2022; Busillo et al. 2023, 2025; Tortora et al. 2025).
@@ Acknowledgements.
VB and CT acknowledge the INAF grant 2022 LEMON. LL thanks the support from INAF theory Grant 2022: Illuminating Dark Matter using Weak Lensing by Cluster Satellites. This research uses observations made with the NASA/ESA Hubble Space Telescope obtained from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555. This work has made use of the Early Release Observations (ERO) data from the Euclid mission of the European Space Agency (ESA), 2024, https://doi.org/10.57780/esa-qmocze3. This work has made use of the Euclid Quick Release Q1 data from the Euclid mission of the European Space Agency (ESA), 2025, https://doi.org/10.57780/esa-2853f3b. The Euclid Consortium acknowledges the European Space Agency and a number of agencies and institutes that have supported the development of Euclid, in particular the Agenzia Spaziale Italiana, the Austrian Forschungsförderungsgesellschaft funded through BMK, the Belgian Science Policy, the Canadian Euclid Consortium, the Deutsches Zentrum für Luft- und Raumfahrt, the DTU Space and the Niels Bohr Institute in Denmark, the French Centre National d’Etudes Spatiales, the Fundação para a Ciência e a Tecnologia, the Hungarian Academy of Sciences, the Ministerio de Ciencia, Innovación y Universidades, the National Aeronautics and Space Administration, the National Astronomical Observatory of Japan, the Netherlandse Onderzoekschool Voor Astronomie, the Norwegian Space Agency, the Research Council of Finland, the Romanian Space Agency, the State Secretariat for Education, Research, and Innovation (SERI) at the Swiss Space Office (SSO), and the United Kingdom Space Agency. A complete and detailed list is available on the Euclid web site (www.euclid-ec.org).
Appendix A Impact of lack of companions in training set on LEMON performance
Figure 14: Same as Fig. 4, but by using LEMON trained only on non companion lenses and using lenses with companions as the test set.
To further check how the completeness of the training set influences both the predictions and the uncertainties of LEMON, we also trained the algorithm only on the lenses without companions. We split the lenses into 40 0004000040\,00040 000 lenses for the training set, $5000$ lenses for the validation set, and $5000$ lenses for the test set. We then trained LEMON with the same procedure as described in Sect. 4.1, and then tested on a companions test set formed by $5000$ randomly extracted lenses from the 50 0005000050\,00050 000 lenses with companions.
The recovery from the model trained without companions in the training set of the parameters for the companions test set is shown in Fig. 14. From the figure, we can see that the Einstein radius is completely unrecovered, with a very strong bias towards higher predicted radii ($\mu\approx 1\arcsec$) and a very low value of the coefficient of determination, $R^{2}=-14.81$, indicating a poor fit. This suggests that the presence of companions heavily affects the recovery of the Einstein radius, potentially because the contaminants are confused with arc features, or because gravitational effects from the companions alter the lensing characteristics of the images, or both. The axis ratio shows a slightly positive value of $R^{2}$, and a systematic underestimate of the predictions, as seen from the metrics, with $\mu=-0.08$ and $\sigma^{-}=0.14$. This indicates that contaminants have a slight effect on the recovery of the lens shape, likely introducing slight distortions of the image that LEMON was not trained to accommodate. The position angle appears to be well recovered, implying that contaminants have practically no effect on the recovery of this parameter. Although the lens effective radius shows a near-zero mean bias, its $R^{2}$ of $-0.47$ and high upper scatter of $0.21″\sigma^{+}=$$$ highlight that a high scatter induced by the light of the contaminants reduces the model’s ability to predict the effective radius, overestimating systematically the size of the galaxy. The lens Sérsic index shows a very large scatter, with $\textrm{NMAD}=1.07$, reflecting LEMON’s difficulty in recovering the light profile slope, potentially due to additional sources of light in the image that confuse the network. Finally, while the magnitude is the best recovered parameter, with a coefficient of determination $R^{2}=0.77$, it is systematically underestimated ($\mu=-0.29$), indicating that the brightness of contaminants causes LEMON to interpret lenses as brighter than they actually are.

### v2

@@ 6.3 LEMON applied to real Euclid Q1 lenses
We finally applied LEMON to the sample of strong lens images found by the strong lensing discovery engine in the Euclid Quick Release 1 (Q1) data (Euclid Collaboration: Walmsley et al. 2025). Through an initial scan of the Q1 galaxy catalogues with machine-learning models, followed by citizen science inspection, expert vetting, and system-by-system modelling, Euclid Collaboration: Walmsley et al. (2025) has collected 500 strong lens candidates, divided in Grade A (246) and B (254) lens candidates after the expert vetting scoring. We complement these data with 38 Grade A and 40 Grade B candidates found after the expert visual classification performed on high-velocity dispersion galaxies in Euclid Collaboration: Rojas et al. (2025).
@@
The strong lenses in Euclid Collaboration: Walmsley et al. (2025) and Euclid Collaboration: Rojas et al. (2025) have been modelled via PyAutoLens, which is the standard algorithm for lens modelling in Euclid  [^10] https://github.com/Jammy2211/PyAutoLens (Nightingale & Dye 2015; Nightingale et al. 2018, 2021b). PyAutoLens is an open-source Python package for strong gravitational lensing, which includes both fully automated strong lens modelling of galaxies and galaxy clusters, and tools for simulating samples of strong lenses. In this work, PyAutoLens employs a SIE mass model, consistent with the model used to train LEMON.
We predicted the mass and light parameters of 354 of these Q1 lenses, filtered such that the classical modelling is successful. Figure 12 shows the comparison between the parameters obtained through the two modelling techniques: from the figure, we can see that the Einstein radius recovery matches the traditional modelling well, with an $R^{2}$ value of $0.71$. The bias and NMAD are very low, respectively, 0.01 and 0.07. The mass ellipticity components do not correlate with PyAutoLens’ modelling results, with all the predictions clustering around zero, while the light ellipticity components match well the classical modelling, with an $R^{2}$ of $0.74\textrm{--}0.78$ and very low values of bias and NMAD. The predicted effective radii from LEMON match those from PyAutoLens up until 1, after which we systematically underpredict with respect to the classical modelling results, similarly to the comparisons with the literature for the effective radius from Sect. 6.1. The Srsic indices predicted by LEMON do not correlate with PyAutoLens’ results. It should first be noted that PyAutoLens’ modelling of these lenses assumes a maximum value for the Srsic index $n=5$. Moreover, we have verified that almost all of the points above 1 in Fig. 12f correspond to points that lie on the $n=5$ strip in Fig. 12g. This could imply that the assumption of a maximum value for the Srsic index biases the classical modelling predictions, leading to differing values with respect to what LEMON predicts. Finally, the lens $I_{\scriptscriptstyle\rm E}$ magnitude has a very good match with classical modelling, having $R^{2}=0.90$. Overall, these results show that LEMON matches well the results of classical modelling for real Euclid lenses, at least for Einstein radius, light ellipticity, effective radius (up to 1) and $I_{\scriptscriptstyle\rm E}$ magnitude (after a zero-point correction).
7 Enhancement of the Euclid pipeline due to LEMON
@@
One possible use of LEMON in the context of strong lensing applications within Euclid, beside fast modelling of strong gravitational lenses by itself, is to both speed-up the modelling of strong gravitational lenses through the Euclid pipeline by giving traditional lens modelling methods initial starting points from which to start the search, and to reduce the failure rate of the modelling pipeline by cross-checking LEMON results with those from the traditional methods.
The Euclid modelling pipeline (Euclid Collaboration: Walmsley et al. 2025) adopts the Python library PyAutoFit [^11] https://github.com/rhayes777/PyAutoFit(Nightingale et al. 2021a) to produce an initial lens model result, which is then used to initialise more complex lens models in the deeper parts of the pipeline (e.g. models that use a Delaunay mesh to reconstruct the source galaxy). PyAutoFit uses nested-sampling algorithms, such as nautilus (Lange 2023), for Bayesian inference of the lens model. The posteriors obtained with this approach, however, are not used for any scientific result; rather, the point-wise results are used as initialisers for the more complex models discussed above.
@@
To efficiently use the information from LEMON, PyAutoLens can also use a gradient ascent optimiser, which fully exploits the information from the initial starting point (Amorisco et al. 2022). The lens light and source galaxy during lens modelling are modelled using a multi-Gaussian expansion technique (see He et al. 2024).
To assess the speed gain on the initial stages of the Euclid lens modelling pipeline, we considered 20 random lenses from the test sample of Euclid mock lenses, and first modelled them through PyAutoLens with the standard modelling pipeline. The modelling took approximately 25 00025\,000 likelihood evaluations, which amount to around $7\,\textrm{h}$ on a single CPU. A total of 16 out of 20 fits successfully recover the correct lens model, demonstrating the reliability of the PyAutoLens standard modelling pipeline. In two of the four bad cases, the fit is biased due to additional line-of-sight emission from a galaxy not associated with the strong lens (a future iteration of the PyAutoLens pipeline will include modelling emission from line-of-sight galaxies).
We then used PyAutoLens with the gradient ascent optimiser, by modelling the lenses with LEMON and then using the obtained best values as starting points for the optimiser. The fits take about $1000$ likelihood evaluations, which amounts to around $0.27\,\textrm{h}$ on a single CPU, which is a substantial gain over the nested-sampling algorithm, with the hybrid method being $26$ times faster. The majority of the fits are good, except for four of the lenses, which return an incorrect lens model. This result aligns with the number of correctly recovered lenses from the standard pipeline alone. It should be noted that, except for one of the lenses having additional line-of-sight emission from a foreground galaxy, which is wrongly modelled by both techniques, the other three lenses not modelled correctly by the hybrid method are different lenses than the ones wrongly modelled by the standard modelling pipeline. This shows another important benefit of combining LEMON with PyAutoLens, namely the ability to complement each other’s limitations: models wrongly inferred by the standard pipeline alone could be modelled correctly by this hybrid approach. Running independent fits with both the standard PyAutoLens pipeline and the combined LEMON-PyAutoLens approach for the same lens, and selecting the best-fit model from the two, will therefore provide a viable strategy to minimise failure rates in lens modelling.
Given that the initialisation stage takes the majority of the time of the pipeline computing time, the speed-up in the initial phase should provide a speed-up of the entire pipeline of up to $5$ times the speed associated with using the nested-sampling algorithm only as an initialiser. This speed-up of modelling is consistent with results from Pearson et al. (2021), where a BNN was used to predict gravitational lens parameters, which are then given to PyAutoLens as priors. In their case, the combination of BNN and PyAutoLens increases the modelling speed by a mean factor of $1.73$. It should be noted, however, that in our case the standard modelling method is not a Markov chain Monte Carlo with priors set by a BNN like in Pearson et al. (2021), and as such, the speed-up factor in our case can be bigger due to the use of a simpler algorithm. This is an optimistic estimate of the speed-up, however, given that on real images the predictions from LEMON are worse than on simulated lenses. In future iterations of LEMON, we plan to implement error estimation on the lens model parameters predicted by PyAutoLens, which are currently not computed with the gradient-ascent method. This can be achieved in different ways, for example by using Hessian matrix estimation or the Fisher information matrix. Investigating these approaches is, however, beyond the scope of this paper.
@@
Finally, to assess whether the optimiser approach needs LEMON in order to correctly model gravitational lenses, we repeated the test by using random starting points for the optimiser. Results show that, with random initialisations, PyAutoLens is unable to successfully recover a correct lens model fit for any of the lenses. Thus, the robust initial starting points provided by LEMON prove to be required in order for PyAutoLens to be sped up by using the gradient ascent optimiser instead of the nested-sampling algorithm.
@@ 8 Conclusions
In this work, we expand the lens modelling machine-learning algorithm LEMON (Gentile et al. 2023) by predicting both the mass and the light of the foreground deflector of Euclid galaxy-galaxy gravitational lenses. The former is modelled as a SIE with external shear, while the latter is a Srsic profile. We used mock Euclid lenses with contaminants in the image for both training and testing of the architecture. The main results are as follows.
@@
  • All parameters of the test set are recovered correctly, with very low values of bias and scatter. The parameter that is recovered worst is the Srsic index, which shows a large scatter and a deviation from the linear trend for high values of the Srsic index. This suggests that LEMON is able to recover the mass and light parameters for the lens even in the presence of companions contaminating the image.
  • Applied to a selection of Euclidised HST images, LEMON manages to match classical modelling in recovering the Einstein radius, mass and light ellipticity components, lens effective radius of the lens up to 1, and lens magnitude for the SLACS sample, while we cannot reach any conclusion about the Srsic index due to low statistics. In general, the results are worse than those obtained from the test set, given that Euclidised lenses may not be similar to the simulations used to train the network, as is shown in Fig. 9.
@@
  • Applied to five real galaxy-galaxy Euclid gravitational lenses observed in the Perseus ERO field (Acevedo Barroso et al. 2025), for which the value of the Einstein radius is available thanks to classical modelling, and the 578 new candidates in the Q1 data (Euclid Collaboration: Walmsley et al. 2025; Euclid Collaboration: Rojas et al. 2025), for which more parameters are available, we match PyAutoLens in recovering the Einstein radius, light ellipticity components, and lens $I_{\scriptscriptstyle\rm E}$ magnitude parameters. The effective radius predictions match PyAutoLens well up to 1, while the Srsic index and mass ellipticities do not correlate as well with the predictions from PyAutoLens. On average, however, the inferred mass ellipticities are not strongly discrepant from the PyAutoLens values.
@@
  • By using the predictions of LEMON as starting points for a gradient-ascent version of the PyAutoLens modelling algorithm, it is possible to speed up the algorithm by up to 26 times compared to the standard Euclid modelling pipeline. This would not be possible without the initial LEMON guesses, because the gradient-ascent optimiser with random initialisations does not produce valid model results for all lenses.
@@
In future works, we plan to further advance the LEMON architecture, by predicting the parameters of more general lens mass models, such as the power-law model. We also plan on enhancing the prediction of certain parameters by deblending the gravitational lens image in lens-only and source-only images (Zhong et al. 2025), and training LEMON separately on the two components.
@@
Q1 data are now public, with DR1 data around the corner. We are therefore preparing to process the enormous amount of upcoming data, and have started applying LEMON to newly discovered strong lenses from the Q1 data release. Results confirm that LEMON is a robust and fast lens modelling tool for Euclid. This will enable the generation of models for the approximately 100 000100\,000 expected Euclid lenses in a remarkably short time.
@@
By applying LEMON and PyAutoLens to Euclid gravitational lenses, we shall determine the total mass within the Einstein radius, or extrapolated to the effective radius, when lens and source redshifts are available. Combining this with stellar mass estimates from SED fitting (Euclid Collaboration: Enia et al. 2025), we shall derive dark matter fractions. We shall also provide constraints on the mass density slope and initial mass function. With such a vast statistical lens sample, we shall be able to investigate the evolution of these properties in unprecedented detail, as well as the scaling relations that relate them as a function of redshift (Koopmans et al. 2006; Gavazzi et al. 2007; Bolton et al. 2008; Auger et al. 2010; Tortora et al. 2010, 2014, 2018, 2019; Sonnenfeld et al. 2013, 2015, 2017; O’Riordan et al. 2025). By comparing these scaling relations with predictions from cosmological simulations, we shall be able to constrain cosmological parameters and physical processes in galaxies (Mukherjee et al. 2018, 2021, 2022; Busillo et al. 2023, 2025; Tortora et al. 2025).
@@ Acknowledgements.
VB and CT acknowledge the INAF grant 2022 LEMON. LL thanks the support from INAF theory Grant 2022: Illuminating Dark Matter using Weak Lensing by Cluster Satellites. This research uses observations made with the NASA/ESA Hubble Space Telescope obtained from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 526555. This work has made use of the Early Release Observations (ERO) data from the Euclid mission of the European Space Agency (ESA), 2024, https://doi.org/10.57780/esa-qmocze3. This work has made use of the Euclid Quick Release Q1 data from the Euclid mission of the European Space Agency (ESA), 2025, https://doi.org/10.57780/esa-2853f3b. The Euclid Consortium acknowledges the European Space Agency and a number of agencies and institutes that have supported the development of Euclid, in particular the Agenzia Spaziale Italiana, the Austrian Forschungsförderungsgesellschaft funded through BMIMI, the Belgian Science Policy, the Canadian Euclid Consortium, the Deutsches Zentrum für Luft- und Raumfahrt, the DTU Space and the Niels Bohr Institute in Denmark, the French Centre National d’Etudes Spatiales, the Fundação para a Ciência e a Tecnologia, the Hungarian Academy of Sciences, the Ministerio de Ciencia, Innovación y Universidades, the National Aeronautics and Space Administration, the National Astronomical Observatory of Japan, the Netherlandse Onderzoekschool Voor Astronomie, the Norwegian Space Agency, the Research Council of Finland, the Romanian Space Agency, the State Secretariat for Education, Research, and Innovation (SERI) at the Swiss Space Office (SSO), and the United Kingdom Space Agency. A complete and detailed list is available on the Euclid web site (www.euclid-ec.org/consortium/community/).
Appendix A Recovery of shear parameters
Figure 13: Same as Fig. 4, but for the external shear components $\gamma_{x}$ and $\gamma_{y}$.
Figure 14: Same as Fig. 12, but for the external shear components $\gamma_{x}$ and $\gamma_{y}$.
As a side test, we also trained LEMON to specifically recover the shear components $\gamma_{x}$ and $\gamma_{y}$, using the second dataset of Euclid mock lenses described in Sect. 2.1 for training, validation and testing. This dataset has the exact same properties as the main dataset, except for the extended interval of values for the shear components.
Figure 13 shows the recovery of the two shear components on the simulated test-set (constructed in an identical way as with the main dataset). This recovery is very good, showing no bias (except for a slight deviation on the edges of the parameter space), low scatter, and a very high value of the coefficient of determination ($R^{2}=0.86\textrm{--}0.87$). Overall, on simulated images LEMON manages to recover these parameters, under the condition that the parameters’ intervals on which the inference is done is larger than the intrinsic scatter on the component. Indeed, when trying to recover the shear components of the main dataset, the inference fails, likely because the parameter values are too close to zero, and the shear effects on the lensing configurations are too small to be discerned. All the other parameters are recovered as well as with the network trained on the main dataset, with similar values for the metrics.
Figure 14 shows a comparison between our estimates of the shear and the estimates from PyAutoLens on the same Q1 lenses used in Sect. 6.3. We can see that the two modelling results strongly disagree, with negative coefficients of determination. This discrepancy is most likely due to the fact that our network finds ambiguity in the mapping from image to shear (hence the very large uncertainties in our predictions with respect to the classical modelling estimates). This ambiguity could be due to several reasons, such as intrinsic degeneracies with mass ellipticity or domain shift issues (see e.g. Filipp et al. 2025).
It should be noted that recovery of external shear parameters with machine learning is notoriously difficult (Gawade et al. 2025; Schuldt et al. 2023a), and as such we did not expect a good match with literature results for real lenses. It is notable, however, that the network manages to recover the external shear values at least in the ideal case of simulated Euclid lenses.